A rail road is moving with a constant acceleration of 1 m/s2 a block of 5 kg is put on a horizontal rough road floor in the car. at time t = 0, velocity of the car is 5m/s . Considering friction is sufficient and block is not slipping on the floor , the work done on the block by friction force during t= 0 to t = 2 sec will be (cofficient of friction is \mu_s\, )???

Grade:11

1 Answers

vikas askiitian expert
13 years ago

seudo force acts on the block placed on car...

Fseudo = ma                            (a is accleration of car)

          =5*1=5N

now since the block is not slipping so friction is sufficient to balance the the suedo force ...

ffriction = Fseudo = 5N

during t=0 to t=2 sec , displacement of car is

S = ut + .5at2

S = 12 m                         (a=1m/s2 , u =5m/s , t=2 )

workdone by friction = f.X = 5*12 = 60N